If you want to be 95​% confident of estimating the population proportion to within a sampling error of plus or minus0.05 and there is historical evidence that the population proportion is approximately 0.46​, what sample size is​ needed?
Answer #1

Solution,

Given that,

\hat p = 0.46

1 - \hat p = 1 - 0.46 = 0.54

margin of error = E = 0.05

At 95% confidence level

\alpha = 1 - 95%

\alpha = 1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96   

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.05 )2 * 0.46 * 0.54

= 381.70

sample size = n = 382